OpenStudy (anonymous):

I am asked to find the equation of the tangent line to the curve f(x)=sin(x)+sin^2(x) at the point (0,0). Can anyone explain how I should approach this problem? I understand that I must differentiate the function, but I can't see how to come up with the equation for the line. Any help would be greatly appreciated.

OpenStudy (zarkon):

can you find the derivative?

OpenStudy (dumbcow):

equation of the line: y = mx+b m = f'(0) **slope is derivative at point (0,0) b = y-mx = 0-f'(0)*0 = 0 --> y = f'(0)x all you need to do is find f'(x)

OpenStudy (anonymous):

slope point form: slope=m point: (x,y) y-y2=m(x-x1) slope=f'[0] point: (0,0) y-0=(f'[x])(x-0)